OpenStudy (anonymous):

I NEED HELP REALLY FAST PLZ!? The system of equations is inconsistent. What are the missing values? y=2/3x+3 y= ?/?x+5

jaynator495 (jaynator495):

now that, thats out of the way

jaynator495 (jaynator495):

oh um this is akward... i cant even help with this... uhh sorry >_<

OpenStudy (anonymous):

Itz ok

OpenStudy (anonymous):

Do u know someone who can help me?

jaynator495 (jaynator495):

@TheSmartOne he should =P

jaynator495 (jaynator495):

or @e.mccormick

TheSmartOne (thesmartone):

2 lines can only be inconsistent when they are parallel.

TheSmartOne (thesmartone):

http://www.purplemath.com/modules/systlin2.htm Look ath that when you have time later :)

TheSmartOne (thesmartone):

ok 2 lines are parallel when they have the same slope but different y-intercepts.

OpenStudy (anonymous):

Ok

TheSmartOne (thesmartone):

the equation is in the form of y=mx+b where m is slope and b is the y-intercept

TheSmartOne (thesmartone):

so what do you think the answer is now? @*louisa* @~ @Jaynator495

TheSmartOne (thesmartone):

And @Jaynator495 was correct when he said that I could help you.

TheSmartOne (thesmartone):

I mean afterall I am TheSmartOne :P

OpenStudy (anonymous):

Lol just a second this is confusing

OpenStudy (anonymous):

I dont get how I would find the answer

OpenStudy (anonymous):

Or where to start to find it

OpenStudy (anonymous):

@TheSmartOne ?

OpenStudy (anonymous):

I have five minutes left on my work

TheSmartOne (thesmartone):

ok.

TheSmartOne (thesmartone):

y=2/3x+3 y= ?/?x+5

OpenStudy (anonymous):

How will I start it

TheSmartOne (thesmartone):

y= ?/?x+5 ^ this has to have the same slope as y=2/3x+3

TheSmartOne (thesmartone):

slope is m in y=mx+b

TheSmartOne (thesmartone):

So both of their m or slopes should equal.

OpenStudy (anonymous):

My sister said of it would be 3/2?

OpenStudy (anonymous):

*if

OpenStudy (anonymous):

Hello?

TheSmartOne (thesmartone):

umm. no

TheSmartOne (thesmartone):